Luzadis Company makes furniture using the latest automated technology. The company uses a job-order costing system and applies manufacturing overhead cost to products on the basis of machine-hours. The predetermined overhead rate was based on a cost formula that estimates $1,520,000 of total manufacturing overhead for an estimated activity level of 76,000 machine-hours. During the year, a large quantity of furniture on the market resulted in cutting back production and a buldup of furniture in the companys warehouse. The companys cost records revealed the following actual cost and operating data for the year 64,000 1,471,000 ufacturing overhead cost Inventories at year-end: Raw materials Work in process (includes overhead applied of $64,000) Finished goods (includes overhead applied of $204,800) 14,000 s 90,500 S 289,600 1,429,900 Cost of goods sold (includes overhead applied of $1,011,200) Required: 1. Compute the underapplied or overapplied overhead. 2. Assume that the company closes any underapplied or overapplied overhead to Cost of Goods Sold. Prepare the appropriate journal entry. (If no entry is required for a transaction/event, select No journal entry required in the first account field.) 3. Assume that the company allocates any underapplied or over appliedoverhead proportionally to Work in Process, Finished Goods, and Cost of Goods Sold. Prepare the appropriate journal entry. (If no entry is required for a transactign/event, select No Journal entry required in the first account field.) 4. How much higher or lower will net operating income be if the underapplied or overapplied overhead is allocated to Work in Process, Finished Goods, and Cost of Goods Sold rather than being closed to Cost of Goods Sold? Complete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 Req 4 Compute the underapplied or overapplied overhead. < Prev 30, 5 ??? Next >
media%2Fa31%2Fa311aebf-9736-4506-af27-30
media%2Fae8%2Fae89a790-7638-484d-86d0-e8
media%2F46f%2F46f032d5-29d3-4b04-8ce9-50
media%2Fac1%2Fac1d3072-9946-4ae2-bc7f-ef
Answer #1
Predetermined overhead rate = 1520000/76000= $20
1
Overhead applied = 64000*20= $1280000
Underapplied overhead = 1471000-1280000= $191000
2
Cost of goods sold 191000
      Manufacturing overhead 191000
3
Cost of goods sold 150890 =191000/1280000*1011200
Work in Process 9550 =191000/1280000*64000
Finished goods 30560 =191000/1280000*204800
      Manufacturing overhead 191000
4
Net operating income will be higher by $40110 (191000-150890)